# SPSysML: A meta-model for quantitative evaluation of Simulation-Physical Systems

### v3

@@ 4 Simulation-Physical System
Simulation-Physical System specification is based on SysML and EARL. In the following sections, we describe SPSysML and SPSysDP. SPSysML defines stereotype-based meta-models of requirements and structure. Description of a meta-model requires a notation for multiple blocks or instances of a stereotype (e.g. two instances of <<Agent>>). For this purpose, we append the stereotype with s (e.g. <<Agent>>s). We refer to an instance of a stereotype using the part-name <<stereotype>> symbol.
4.1 SPSysML – requirements meta-model for SPSys
@@
SPSysML specifies the requirements model as SysML profile. The requirements can be defined based on various premises. In particular, they can derive from user requirements [47]. The SPSysML profile defines structural, functional, configurational and environment requirements. Other requirement types can be added; however, these are used by SPSysDP in system setups’ specification and requirement-based decomposition of the system. Auxiliary sequence diagrams presenting the concept of the system behaviour and use case diagrams are useful for requirements definition. Before SPSys requirements specification, its environment must be analysed. Therefore, in SPSysML, we distinguish Environment requirements. They can be of various stereotypes. This article focuses on <<ExogAgentReq>> that describe exogenous agents interacting with the world alongside the system. In SPSysML, the system is decomposed into parts based on <<SysPartReq>>s. Each <<SysPartReq>> is classified to elementary <<HardwareReq>> or more general <<PhyPartReq>>, <<SimPartReq>> or <<SimPhyPartReq>>. <<PhyPartReq>> determines a part interacting only with the Physical World (even during the system development, e.g. during parallel development of the dependent parts), and its simulated embodiment is not required (or cannot be created) during the system development. <<SimPartReq>> specifies a part interacting only with the Simulated World (e.g. mock-ups or demonstrators). <<SimPhyPartReq>> interacts with both Worlds (e.g. realised with a pair of DT and PT). <<HardwareReq>>s are in a satisfy relationship with <<FunctionalReq>>s. This means hardware requirements specified with a <<HardwareReq>> enable the realisation of the functionality specified with the given <<FunctionalReq>>. The configurational stereotypes specify if a system part is obligatory to launch in any of its configurations (<<ObligatoryReq>>) or if it is not (<<OptionalReq>>). With optional requirements, the system gains multiple setups, addressing the optional requirement or not. The proposed development method defines separate system component groups, each structuring a system setup. The system setups are defined for any combination of the optional requirements.
4.2 SPSysML – structure meta-model for SPSys
SPSysML derives from EARL [37] version 1.3 [48]. We use a Group of Subsystems (<<GpSubsys>>) to gather Subsystems with a specific common properties, and a Group of Agents (<<GpAgents>>) to organise the Agents (composed of Subsystems) cooperating for a defined aim in the system. In a SPSys we differentiate three specific Agent stereotypes that derive from <<Agent>> defined in EARL (Fig. 5):
@@ 4.2 SPSysML – structure meta-model for SPSys
  • Hybrid Agent (<<HybAgent>>) – Runs in both embodiments with neither interaction nor perception of any world.
Figure 5: Simulation-Physical System composition
The Groups of Agents called World Mirror Group of Agents aggregate <<SimAgent>>s that are Digital Twins of the Agents in the Physical World that are not controlled by the system and execute exogenous actions in the Simulated World. Groups of this type exist if the system works in a dynamic environment (e.g. an environment with human inhabitants). For example, a World Mirror Group of Agents modifies the Simulated World as humans do in the Physical World.
@@ 4.2.1 Digital Twins in SPSys
Integrity between SPSys embodiments is crucial. Therefore, the existence of DT of physical parts is advisable and common. We define the mirror relationship between <<PhyAgent>> and <<SimAgent>> to model the relationship between DT and its PT. To enable multi-agent DT for a single <<PhyAgent>> or single-agent DT for multi-agent PT, we introduce <<MirrPhyGpAgents>> and <<MirrSimGpAgents>>. They aggregate Agents of different stereotypes (Fig. 6).
@@
It is worth noting that the same stimuli of the mirroring Groups cause corresponding reactions in specific worlds, in the Simulated World for Mirror Simulation Group of Agents and in the Physical World for Mirror Physical Group of Agents. The particular reaction results from the requirements of the specific system and does not need to be identical between the Agent Groups that mirror each other. A stimuli of a Group of Agents is the input from other Agents or the environment through the Group of Agents interfaces. Agent’s reaction to a stimuli manifests as the Agent state change (including changes in memory or output data). Considering a pan-tilt head as an example, <<MirrSimGpAgents>> is the simulated head, and <<MirrPhyGpAgents>> is the real one. In this case, the common stimuli are the light perception from the according world (simulated/physical) and the head joints’ control. The common reaction is data in their output interface. A digital twin (<<MirrSimGpAgents>>) for a pan-tilt head development is a convenient tool, e.g. for servovision implementation or joint constraints evaluation. SPSys may consist of multiple such pairs, e.g. manipulator and mobile base in a service robot case.
The mirror relationship does not propagate automatically to all the Agents aggregated in the mirroring Groups, so agents from <<MirrSimGpAgents>> do not mirror all agents of mirroring <<MirrPhyGpAgents>> and vice versa. The number of mirroring Agents should be maximised to maximise coverage of simulation-based testing and profits of DT. Additionally, the complexity and quantity of mirroring Agent Subsystems should be minimised to boost the modularity and integrity of the system. Otherwise, the system could be composed of just two mirroring agents. If a computational functionality is required in both embodiments, a <<HybAgent>> should be designed for this purpose.
@@ 4.2.2 Subsystem types
The Agents are built with Subsystems of different types based on EARL. In EARL the central computational part of an Agent is the Control Subsystem <<ContSubsys>>. An Agent communicates with other Agents using communication buffers of its <<ContSubsys>>. To percept the environment, an Agent uses Real Receptors (<<RealRec>>s), and to aggregate and preprocess stimuli, it uses Virtual Receptors (<<VirtRec>>s). To affect the environment, an Agent uses Real Effectors (<<RealEff>>s). To preprocess <<ContSubsys>> commands to signals for <<RealEff>>s it uses Virtual Effectors (<<VirtEff>>s). SPSys interacts with both the simulated and physical world; thus, SPSysML differentiates between the types of the above Subsystems— Simulated, Physical, and Simulated-Physical. To achieve maximum integrity between the simulated and physical embodiments, all <<ContSubsys>>s should be Simulated-Physical (<<SimPhyContSubsys>> stereotype) and constitute the general concept of the shared controller, recall Fig. 1. For iterating design purposes, embodiment-specific Control Subsystems concepts (<<PhyContSubsys>> and <<SimContSubsys>>) may be helpful. The other Subsystems can be Simulated or Physical. SPSysML defines <<GpSubsys>>s that aggregate receptors and effectors considering their embodiment and the world they interact with (Fig. 7, Fig. 8). As one <<SimPhyContSubsys>> can communicate with multiple Simulated/Physical Virtual Receptors and Effectors, we define four<<GpSubsys>>s aggregating Physical/Simulated Virtual/Real Subsystems. The data flow and communication links for <<SimPhyContSubsys>> interacting with Simulated and Physical Worlds is shown in (Fig. 9).
@@
4.2.3 Realisation of the Subsystem and Agent types
[…]
@@ 4.2.3 Realisation of the Subsystem and Agent types
It should be noted that the Basic EARL meta-model defines only one <<ContSubsys>> in an Agent, and inter-agent communication is handled only by the <<ContSubsys>>. In SPSysML, we differentiate between specific types of <<ContSubsys>> for the embodiments; still, an Agent can aggregate just one specialisation of <<ContSubsys>>. SPSys can be deployed in various setups (e.g. testing setups). The setups are a Group of Agents fulfilling the system’s functionality in the setup. For the system setup definition, we use <<SysSetupGpAgents>> to specify the Group of Agents working in the setup. The specific <<SysSetupGpAgents>> is derived from the system’s requirements and test scenarios. This problem is considered in the SPSysDP section (Sec. 4.3).
4.3 Simulation-Physical System Development Procedure (SPSysDP)
SPSysML defines system parts and the relations between them to enable SPSys specification. Various agent-based system development procedures can be used for SPSysML-based systems; however, we define one that enables inter-embodiment integrity optimisation for SPSys (Fig. 11). The main activities of the procedure can be executed in different development approaches, traditional, agile, or hybrid. In agile and hybrid approaches, the system is specified and implemented partially in an iterative manner. In the procedure shown in Fig. 11, there are four decision nodes D1, D2, D3, and D4. The exact development procedure results from the logic predicates set to these nodes. D1 checks if the requirements of the considered system part are comprehensive and if the system parts and functions, considering the system’s test scenarios, are expressed in the requirement diagrams. D2 checks if the structure evaluation is satisfactory. D3 checks whether any detailed analysis or requirements modification is required based on the design iteration, and D4 checks if the required <<SysSetupGpAgents>>s are implemented and tested successfully. If needed, the procedure enables complete reiteration from the design stage.
Figure 11: SPSys Development Procedure for Simulation-Physical Systems
System planning and analysis
  • Step 1 (Requirement specification): The requirements are specified following the model defined in Sec. 4.1. It is recommended to define system use cases and typical interactions between system parts on sequence diagrams while specifying the requirements.
  • Step 2 (Requirement analysis): Analysis of the requirements considering the project stakeholders’ demands and the system’s test scenarios. In case it is needed, modifications to these requirements are made.
System design
  • Step 3 (World Synchronization): This step is not applicable if the system works in a static environment (there are no exogenous actions of any non-system agent executed on the system’s environment). In other cases, simulation of external agents (e.g. humans) is required, and it is done by WorldSync <<WorldMirrGpAgents>>. Each <<SimAgent>> composed in <<WorldMirrGpAgents>> manages one <<ExogAgentReq>> specified in Steps 1-2. The key design aspects for the <<SimAgent>>s are the Simulated World model perceived by the SPSys’ receptors and the actions affecting this World. To model humans as <<SimAgent>>, we propose our framework Human Behaviour in Robotics Research (HuBeRo) [49, 50]. It specifies and implements agents mirroring human behaviours and their physical models in simulation (Fig. 12).
    Figure 12: Agent types of HuBeRo framework [49] used in SPSys as <<WorldMirrGpAgents>> that mirrors humans in Simulated World.
  • Step 4 (System decomposition): The system is decomposed into Agents systematically, based on the parts in the requirements. For each <<PhyPartReq>> and <<SimPartReq>>part <<PhyAgent>> or part <<SimAgent>> are created. Each <<SimPhyPartReq>> that has only <<ComputationalFunReq>> becomes part <<HybAgent>>. The other <<SimPhyPartReq>>s become pairs of DT/PT, thus, are realised with a pair of mirroring part <<MirrPhyGpAgents>> and part <<MirrSimGpAgents>> (Fig. 13),
  • Step 5 (DT/PT decomposition – Mirroring Agent Groups specification): Each mirroring part <<MirrPhyGpAgents>> and part <<MirrSimGpAgents>> is iteratively decomposed to more groups to finally reach mirroring groups that each can be realised with a single <<PhyAgent>> or <<SimAgent>>. Each result of the decomposition iteration represents a layer of the initial Group. The mirror relationship is set between <<MirrSimGpAgents>> and <<MirrPhyGpAgents>> and is specified in each layer of the decomposition. The decomposition scheme is shown in Fig. 13.
    Figure 13: Requirement-based SPSys decomposition, where mirror relationship constituting Physical and Digital Twins is obligatory for <<SimPhyPartReq>>. As a result of the decomposition, DT/PT may share <<HybAgent>>s.
  • Step 6 (Agent decomposition): Each <<Agent>> is decomposed to Subsystems. Various approaches can be used depending on the requirements formulation; however, we advise a layered bottom-up one with a definition of each layer interface. First, the hardware assigned to each <<Agent>> is expressed as <<PhyRealRec>>s, <<PhyRealEff>>s, <<SimRealEff>>s and <<SimRealRec>>s (constituting Agent hardware layer). Next, the interfaces of <<PhyVirtRec>>s, <<PhyVirtEff>>s, <<SimVirtEff>>s and <<SimVirtRec>>s are defined based on Sec. 4.2.2 (constituting Agent driver layer). The target is to develop a <<SimPhyContSubsys>> (constituting Agent control layer) that:
    • manages both mirroring Agents’ behaviour,
    • controls hardware (simulated/physical) using the embodiment-common interface of <<PhyDriverGpSubsys>> and <<SimDriverGpSubsys>>,
    • communicates with other Agents using an inter-agent interface.
    However, during the design process and testing, a temporary <<SimContSubsys>> and <<PhyContSubsys>> are helpful. Mirroring Agents consist of identical control layers; however, the Agents differ in drivers and hardware layers. The driver layer must fill the gap between the hardware and controller layers in the specific embodiments to enable embodiment abstraction for the control layer in mirroring Agents.
  • Step 7 (Structure evaluation): In each design iteration, the structure is evaluated. We propose the following evaluation factors for a two-scope analysis:
@@ 4.3 Simulation-Physical System Development Procedure (SPSysDP)
[…]
    The structure optimisation target is the maximisation of the embodiment-common part of the system (IIF, MIF, DGF, DTC factors) to maximise simulation-based testing coverage. The share of physical parts mirrored with Digital Twins (DTC) additionally boosts the system’s robustness, as DT can replace malfunctioned hardware. The detailed analysis of the proposed factors is described in Sec. 5. The optimisation target can be defined with the above factors or others that can be defined for a specific system. If the evaluation result is unsatisfactory, the system should be redesigned (starting from Step 5) following the factors’ maximisation advice defined in Sec. 5.
  • Step 8 (System setups): Based on the optional stereotypes in the requirements, all possible system setups emerge. Each setup (<<SysSetupGpAgents>>) is specified with an Internal Block Diagram (example diagrams are shown in Fig. 1819). The diagram shows the Agents composing given <<SysSetupGpAgents>>, their communication and interaction with Simulated and Physical Worlds. It should be noted that the system’s operational setups are just a starting point for <<SysSetupGpAgents>>s specification. For each implementation testing scenario, a <<SysSetupGpAgents>> should be specified. DTs are broadly used in development optimisation and CPS development safety improvement; therefore, in particular, for testing a part <<MirrPhyGpAgents>>, a <<SysSetupGpAgents>> consisting of part <<MirrSimGpAgents>> (DT of the <<MirrPhyGpAgents>>) should be defined. The next step of this procedure makes sure that the <<SysSetupGpAgents>> with part <<MirrSimGpAgents>> is implemented and tested prior to <<SysSetupGpAgents>> including part <<MirrPhyGpAgents>>.
System implementation and testing
  • Step 9 (Implementation of the Agents): <<SimAgent>>s and <<HybAgent>>s should be implemented before <<PhyAgent>>s for safety reasons. We want to test the system in simulation before its deployment on the real hardware. To achieve this, we queue <<SysSetupGpAgents>>s implementation and testing based on the simulation implementation factor, such that we begin with a <<SysSetupGpAgents>> consisting of the highest share of the <<SimAgent>>s and <<HybAgent>>s and the lowest of <<PhyAgent>>s. The share evaluation is achieved with the simulation implementation factor. Therefore, the implementation step starts with a <<SysSetupGpAgents>>, whose simulation implementation factor is the highest. The simulation implementation factor for example ex1 <<SysSetupGpAgents>> is $\mathtt{SIF}_{ex1}=\frac{s_{ex1}+h_{ex1}}{a_{all}}$, where $s_{ex1}$ and $h_{ex1}$ are cardinalities of the unimplemented <<SimAgent>>s and <<HybAgent>>s in ex1 <<SysSetupGpAgents>> and ${a_{all}}$ is a number of Agents in the system. Before implementation, Subsystems must be translated to a PSM. For robotic systems utilising ROS/ROS2, we propose MeROS DSL [51]. Implementation should include unit testing of each Subsystem and Agent (e.g. applying Robot Unit Testing methodology [30]).
  • Step 10 (Integration verification): Test scenarios execution for all fully implemented <<SysSetupGpAgents>>.
5 The design evaluation factors analysis
This section describes the interpretation of the evaluation factors and which system features they evaluate. We describe characteristics of edge case SPSys that scores maximum or minimum values of the design evaluation factors. This gives a basic intuition on the correlation between the factor values and the SPSys features. The interpretation of the evaluation factors is as follows:
  • IIF – is a share of the software controller common between the embodiments. Virtual Subsystems are not considered, as their count may be related to the Real Subsystem counts in each embodiment. It is maximised by the reduction of the number of <<SimContSubsys>>s and <<PhyContSubsys>>s in favour of an inter-embodiment <<SimPhyContSubsys>>s, The higher IIF is, the more software components are shared between the simulation and physical embodiments. At maximum (IIF=1), all hardware abstract parts of the system are common.
@@ 5 The design evaluation factors analysis
[…]
@@
Based on the factors’ values, one can evaluate the system design in terms of the following:
[…]
@@
Maximisation of the factors is not always required, and the optimisation goal can be set at a different point in the factors’ space. The goal depends on the specific system requirements. However, the factors’ values inform the designer about the inter-embodiment integrity of the system design, so her/his decision is conscious.
6 Ilustrative example - goal validation
We execute the SPSysDP to design, implement and test a complex SPSys utilising a service robot for the INCARE project. The INCARE system idea and requirements are published in [52]. The framework model developed to manage robot tasks is published in [53, 54]. Developing a complex system requires different implementations of its parts in various development phases. In INCARE, we use some commercial products like the TIAGo robot with its control system. We develop and integrate new parts into it (e.g. human fall detector and TIAGo audio interface extension). In such a case, developing one part requires a dummy of another part while being developed simultaneously. Developers can simulate the underdeveloped parts while the physical devices are under construction or development. In the case of INCARE, we use a dummy for the human fall detector while developing a TIAGo robot application helping the elderly who may fall over. We use and modify the TIAGo robot in this project; thus, we specify its hardware and controller as a part of the INCARE system specification. The result of each step of SPSysDP is as follows:
Step 1 and 2
we specify the structural and functional requirements based on the general requirements of the INCARE project. The general and the robot-specific requirement diagrams are shown in Fig. 14.

### v5

@@ 4 Simulated-Physical System
Simulated-Physical Systems Modeling Language (SPSysML) is based on SysML and EARL. In the following sections, we describe SPSysML, which defines a stereotype-based meta-model of requirements and system structure. Description of a DSL requires a notation for multiple blocks or instances of a stereotype (e.g. two instances of <<Agent>>). For this purpose, we append the stereotype with s (e.g. <<Agent>>s). We refer to an instance of a stereotype using the part-name <<stereotype>> symbol.
4.1 SPSysML – requirements meta-model
@@
SPSysML specifies the requirements model as SysML profile. The requirements can be defined based on various premises. In particular, they can derive from user requirements [51]. The SPSysML profile defines structural, functional, configurational and environment requirements. Other requirement types can be added; however, these are used in system setups’ specification and requirement-based system composition. Auxiliary sequence diagrams presenting the concept of the system behaviour and use case diagrams are useful for defining requirements. Before SPSys requirements specification, its environment must be analysed. In SPSysML, we distinguish <<ExogAgentReq>> being a part of the environment requirements that specify exogenous agents interacting with the world alongside the system. The system functions given by <<FunctionalReq>> stereotyped requirements are classified into:
  • simulated-physical, which require perceiving or affecting the simulated/physical world— <<SimPhyFunReq>>,
  • computational, which do not interact directly with any simulated/physical world— <<CompFunReq>>.
Each <<FunctionalReq>> requirement must be performed by a system part; thus <<SysPartReq>> stereotyped requirements must be specified. As shown in Fig. 3, they must satisfy, verify, and derive from their origin— <<FunctionalReq>>. In SPSysML, the system structure is based on the requirements of <<SysPartReq>>s stereotype. Each <<SysPartReq>> is classified as elementary <<HardwareReq>> or one of more general <<PhyPartReq>>, <<SimPartReq>> or <<SimPhyPartReq>>.
  • <<PhyPartReq>> determines a part interacting only with the Physical World (even during the system development, e.g. during parallel development of the connected parts), and its simulated embodiment is not required (or cannot be created) during the system development.
  • <<SimPartReq>> specifies a part interacting only with the Simulated World (e.g. mock-ups or demonstrators). <<SimPhyPartReq>> interacts with both Worlds (e.g. realised with a pair of DT and PT).
  • <<HardwareReq>>s are in a satisfy relationship with <<FunctionalReq>>s. This means hardware requirements specified with a <<HardwareReq>> enable the realisation of the functionality specified with the given <<FunctionalReq>>.
The configurational stereotypes specify if a system part (<<SysPartReq>>) or a system function (<<FunctionalReq>>) is obligatory to launch the system in all of its setups. Configurational stereotypes are orthogonal to structural and functional stereotypes; thus, functional and structural requirements may be either <<ObligatoryReq>> or <<OptionalReq>>. However, a <<SysPartReq>> realising a <<FunctionalReq>> derives its configurational stereotype from the functional requirement. Each <<OptionalReq,FunctionalReq>> gives the system more setups, and the number of the setups equals $2^{n}$, where $n$ is the number of <<OptionalReq,FunctionalReq>>. As a result,
  • <<ObligatoryReq,FunctionalReq>> requirements define the core system’s functions,
  • <<OptionalReq,FunctionalReq>> requirements manage the system’s versatility by dividing the system into system setups with different functionality,
  • <<ObligatoryReq,SysPartReq>> requirements define the system’s core components,
  • <<OptionalReq,SysPartReq>> requirements define the system’s components for different functionality setups.
4.2 SPSysML – structure of SPSys
SPSysML derives from EARL [31] version 1.3 [52]. In a SPSys we differentiate three specialisations of the <<Agent>> stereotype defined in EARL (Fig. 4):
@@ 4.2 SPSysML – structure of SPSys
  • Hybrid Agent (<<HybAgent>>) – Runs in both embodiments and computes without interacting or perceiving any world.
Figure 4: Simulated-Physical System composition
The Groups of Agents called World Mirror Group of Agents aggregate <<SimAgent>>s that are Digital Twins of the Agents in the Physical World that are not controlled by the physical embodiment of the system. They execute exogenous actions in the Simulated World. Groups of this type exist if the system works in a dynamic environment (e.g. an environment with human inhabitants). For example, a World Mirror Group of Agents modifies the Simulated World as humans do in the Physical World.
An agent is composed of Subsystems of different types as specified in EARL. Types of Subsystems are defined in 4.2.2. We use a Group of Subsystems (<<GpSubsys>>) to gather Subsystems with a specific common properties, and a Group of Agents (<<GpAgents>>) to organise the Agents (composed of Subsystems) cooperating for a defined aim in the system.
@@ 4.2.1 Digital Twins in SPSys
Integrity between SPSys embodiments is crucial. Therefore, creating DT of a physical parts is advisable and common. We define the mirror relationship between <<PhyAgent>> and <<SimAgent>> to model the relationship between DT and its PT. To enable multi-agent DT for a single <<PhyAgent>> and vice versa, we introduce <<MirrPhyGpAgents>> and <<MirrSimGpAgents>>. They aggregate Agents of different stereotypes (Fig. 5).
@@
It is worth noting that the same stimuli of the mirroring Groups cause corresponding reactions in specific worlds, in the Simulated World for Mirror Simulation Group of Agents and in the Physical World for Mirror Physical Group of Agents. The particular reaction results from the requirements of the specific system and does not need to be identical between the Agent Groups that mirror each other. A stimuli for a Group of Agents is the input from other Agents or the environment through the interfaces. Agent’s reaction to a stimuli manifests as the Agent state change (including changes in memory or output data).
Example DT-PT pair
Considering a pan-tilt head as an example, <<MirrSimGpAgents>> is the simulated head, and <<MirrPhyGpAgents>> is the real one. The stimuli of this DT-PT pair include the light perception from the according world (simulated/physical) and the control for the head joints. The common reaction is data in their output interface. The DT (<<MirrSimGpAgents>>) for a pan-tilt head development is a convenient tool, e.g. for servovision implementation or joint constraints evaluation. SPSys may consist of multiple such pairs, e.g. manipulator and mobile base in a service robot case.
Constaints and guides for DT-PT pair
The mirror relationship does not propagate automatically to all the Agents aggregated in the mirroring Groups, so agents from <<MirrSimGpAgents>> do not mirror all agents of mirroring <<MirrPhyGpAgents>> and vice versa.
The number of Agents outside DT/PT pairs should be minimised to maximise coverage of simulation-based testing and analysis. Additionally, the complexity and quantity of mirroring Agent Subsystems should be minimised to boost the modularity and integrity of the system. Otherwise, the system could be composed of just two mirroring agents. If a computational functionality is required in both embodiments, a <<HybAgent>> should be designed for this purpose.
@@ 4.2.2 Subsystem types
The Agents are built with Subsystems of different types. In EARL, the central computational part of an Agent is the Control Subsystem <<ContSubsys>>. An Agent communicates with other Agents using communication buffers of its <<ContSubsys>>. To percept the environment, an Agent uses Real Receptors (<<RealRec>>s), and to aggregate and preprocess stimuli, it uses Virtual Receptors (<<VirtRec>>s). To affect the environment, an Agent uses Real Effectors (<<RealEff>>s). To preprocess <<ContSubsys>> commands to signals for <<RealEff>>s, it uses Virtual Effectors (<<VirtEff>>s). Detailed description and SysML diagrams of Agent structure and behaviour are published in [31].
SPSys interacts with both the simulated and physical world; thus, SPSysML differentiates between the types of the above Subsystems— Simulated, Physical, and Simulated-Physical. To achieve maximum integrity between the simulated and physical embodiments, all <<ContSubsys>>s should be Simulated-Physical (<<SimPhyContSubsys>> stereotype) and constitute the general concept of the shared controller, recall Fig. 1. For iterating design purposes, embodiment-specific Control Subsystems concepts (<<PhyContSubsys>> and <<SimContSubsys>>) may be helpful. Virtual and real effectors/receptors can be Simulated or Physical. In SPSysML we group them in four <<GpSubsys>> that aggregate receptors and effectors for each embodiment— simulated and physical (Fig. 6, Fig. 7). Real effectors/receptors are hardware parts, and Virtual effectors/receptors are their drivers.
In complex systems, one <<SimPhyContSubsys>> may communicate with many Simulated/Physical Virtual Receptors and Effectors. In this case, the <<SimPhyContSubsys>> is reused in <<SimAgent>> and <<PhyAgent>>, and both utilise embodiment-specific hardware and drivers. The data flow and communication links for <<SimPhyContSubsys>> interacting with Simulated and Physical Worlds are shown in (Fig. 8).
@@
Agent composition constraints
Each Agent type defined in SPSysML aggregates a particular number of Subsystems of a specific type (Fig. 9).
@@
It should be noted that the Basic EARL meta-model defines only one <<ContSubsys>> in an Agent, and inter-agent communication is handled only by the <<ContSubsys>>. In SPSysML, we differentiate between specific types of <<ContSubsys>> for the embodiments; still, an Agent can aggregate just one specialisation of <<ContSubsys>>.
System setups composition
SPSys can be deployed in various setups (e.g. testing setups). The setups are a Group of Agents fulfilling the system’s functionality in the setup. For the system setup definition, we use <<SysSetupGpAgents>> to specify the Group of Agents working in the setup. The set of agents composing <<SysSetupGpAgents>> is derived from the system’s requirements and test scenarios.
Example Physical embodiment realisation
Subsystems in a <<PhyHardGpSubsys>> (sensors/actuators) communicate with their drivers via various interfaces (e.g. Linux kernel driver for Inter-Integrated Circuit (I2C) or network interface). <<PhyDriverGpSubsys>>s take data from these interfaces, aggregate it and expose their interface for a <<ContSubsys>> executing the logic of the Agent.
Example Simulated embodiment realisation
Subsystems in <<SimHardGpSubsys>> expose Application Programming Interface (API) classes, defined in the simulation environment (e.g. gazebo::ModelPlugin for <<SimRealEff>> and gazebo::SensorPlugin for <<SimRealRec>> in Gazebo444a popular simulation environment used in the DARPA Robotics Challenge, in July of 2013). Subsystems of <<SimDriverGpSubsys>> connect to these interfaces using the API, aggregate sensor data, and respond <<ContSubsys>>s control.
The connections between the groups are shown in Fig. 8.
5 The design evaluation factors
We recognise the following integrity factors classified for design analysis in two scopes:
@@ 5 The design evaluation factors
    • Mirror integrity factor (MIFn=$\frac{c_{n}^{dcs}}{c_{n}^{All}}$), where n is the considered pair of mirroring <<MirrPhyGpAgents>> and <<MirrSimGpAgents>> composing one DT/PT pair, $c_{n}^{dcs}$ and $c_{n}^{All}$ are the counts of <<SimPhyContSubsys>>s and all <<ContSubsys>>s accordingly in nth DT.
Below, we present the interpretation of the evaluation factors and the correlation between the factor values and the SPSys features. We do it by describing the characteristics of edge case SPSys that scores maximum or minimum values of the design evaluation factors:
  • IIF – is a share of the software controller common between the embodiments. Virtual Subsystems are not considered, as their count may be related to the Real Subsystem counts in each embodiment. It is maximised by the reduction of the number of <<SimContSubsys>>s and <<PhyContSubsys>>s in favour of an inter-embodiment <<SimPhyContSubsys>>s, The higher IIF is, the more software components are shared between the simulated and physical embodiments. At maximum (IIF=1), all hardware abstract parts of the system are common.
@@
  • MIFn – is a share of the system parts common between Physical and Digital Twins composing the nth twin pair (managing given <<SimPhyPartReq>>). It is similar to IIF but within the scope of $n^{th}$ pair of Physical and Digital Twin. Tips for maximisation of the MIFn factor are:
@@
Based on the factors’ values, one can evaluate the system design in terms of:
  • Safety and hardware independence during software testing – based on IIF (for system scope), MIFn (for $n^{th}$ DT),
@@
Maximisation of these factors is not always required, and the optimisation goal can be set at a different point in the factors’ space. The goal depends on the specific system requirements. However, the factors’ values inform the designer about the inter-embodiment integrity of the system design, so her/his decision is conscious.
6 SPSysML Validation
First, we describe the SPSysML validation by application. It includes the requirement-based system composition, design evaluation, and applicability in the system development procedure 6.1. Next, we share the results of the SPSysML assessment done among systems engineering practitioners, including a third-party SPSysML user 6.2.
6.1 Example system design and analysis
SPSysML was used in developing complex SPSys utilising a service robot for the INCARE project. This is the primary validation of SPSysML. The INCARE system idea and requirements are published in [53]. The framework model developed to manage robot tasks is published in [54, 55]. Developing a complex system requires different implementations of its parts in various development phases. In INCARE, we use some commercial products like the TIAGo robot with its control system. We develop and integrate new parts into the system (e.g. human fall detector and TIAGo audio interface extension). We use and modify the TIAGo robot in this project; thus, we specify its hardware and controller as a part of the INCARE system specification. We present the result of the SPSysML application in the following SPSys development tasks.
Requirement engineering
In these steps, we use the requirement model defined in SPSysML. We specify the structural and functional requirements based on the general requirements of the INCARE project. The general and the robot-specific requirement diagrams are shown in Fig. 10.